# Tag-Instruct: Controlled Instruction Complexity Enhancement through Structure-based Augmentation

### v1

@@ Abstract
Tag-Instruct: Controlled Instruction Complexity Enhancement through Structure-based Augmentation
[…]
@@ 1 Introduction
Figure 1: AlpacaEval 2 Length Control Win Rate. Tag-Instruct achieves the highest win rate among all compared methods. Blue, yellow, and gray bars represent LLaMA3-8B trained with Qwen2.5-7B-Instruct, Mistral-8B-Instruct, and Alpaca-5k respectively. Evaluated using GPT-4o-Mini.
While recent work has made progress in synthetic data generation Zhu et al. (2024); Xu et al. (2024), current instruction datasets still lack sufficient complexity to fully develop model capabilities Xu et al. (2023). Studies show that model performance strongly correlates with instruction difficulty Zhao et al. (2023, 2024), yet existing approaches struggle to effectively control and scale instruction difficulty. Specifically, Xu et al. (2023) and its follow-up works Wang et al. (2024b); Zeng et al. (2024) rely on models’ prior knowledge for self-reflection on augmentation directions, typically using prompts to directly generate “harder” instructions. However, this approach lacks precise control over difficulty progression, as the concept of “harder” remains ambiguous and unquantifiable, making it challenging to systematically guide the augmentation process. While Zhao et al. (2023) attempts to address this by discretizing instructions into semantic trees for manipulation, it still struggles to identify which semantic components are worth augmenting and how to effectively guide the augmentation direction [^1] Even the implementation of tree-instruct is essentially prompt-based. We name these approaches as Prompt-based Augmentation, which face two fundamental challenges: (1) difficulty in identifying which semantic components are worth augmenting, and (2) lack of principled guidance on which augmentation direction would lead to meaningful complexity increases.
@@
Building on this intuition, we propose Tag-Instruct, a novel Compress & Operation framework for controlled difficulty augmentation. Our framework operates through a three-stage iterative pipeline: (1) semantic encoding that compresses instructions into a compact tag representation while preserving task-relevant information, (2) RL-based tag expansion that explores the tag space to discover new semantic tags that meaningfully increase complexity.
(3) instruction synthesis that generates enhanced instructions by conditioning on both the expanded tag set and original instruction. This iterative process enables progressive difficulty increases while maintaining semantic coherence.
Extensive experiments demonstrate the effectiveness of Tag-Instruct across diverse settings. Using Alpaca-5k Taori et al. (2023) as the base dataset, our method enables LLaMA-3-8B to achieve performance comparable to GPT-4-Turbo on standard benchmarks, significantly outperforming existing instruction augmentation approaches. Through detailed ablation studies, we verified that operating in tag space provides superior controllability and makes different instruction synthesis frameworks more stable and effective.
@@ 3.1 Motivation and Overview
The quality of synthesized data correlates to the correctness, diversity and complexity Liu et al. (2023b). Previous works focus on enhancing the diversity of synthesized instruction data while paying less attention to data complexity. We propose Tag-Instruct to further optimize the complexity of the instruction data synthesized with various frameworks like self-instruct Wang et al. (2023) and evol-instruct Xu et al. (2023) with the help of a teacher LLM. In contrast to prior approaches which directly manipulate raw text through prompting LLMs, Tag-Instruct first compresses the instructions into a low-dimensional tag space (Section 3.2) and then enhances the tag complexity through tag expansion (Section 3.3) guided with tag utility estimation (Section 3.4). At last, the LLM is prompted to generate the optimized instruction based on the original instruction and the enhanced tags (Section 3.5). Following the practice of earlier works, we focus on the optimization of input instructions while the output responses are directly generated by prompting the teacher LLM. The Tag-Instruct works in an iterative manner, starting from each instruction in the initial data, the instruction are optimized through multiple iterations with the Tag-Instruct framework. Please refer to Figure 2 for the overview of Tag-Instruct framework.
@@ 3.2 Semantic Compression
Given an instruction $x$ from existing synthesized data, the teacher LLM is prompted (see the instruction-to-tag prompt in Appendix H) to generate relevant tags $z_{\mathrm{base}}$ of the input instruction. The crafted prompt is expected to explicitly guide the model to extract tags that preserve core semantic meanings with the input. Among these tags, those related to action-oriented concepts and task objectives are emphasized and with higher priority as they are indicative of the input intension. For example, the instruction Create an HTML page that displays a list of books and their authors is compressed into {book_listing, author_info, web_develop}. We discuss and compare different prompt guidelines for semantic compression in Section 5.1.
@@ 3.3 Tag Complexity Expansion
While semantic alignment and global coherence can be effectively verified through prompt-based method, we focus on quantifying the tag utility. A naive approach would be directly prompting the LLM to evaluate the tag utility Zheng et al. (2023). However, the prompt-based method may not fully capture the complex interactions between tags, as shown in our analysis in Section 5.2.
Tag-Instruct is proposed to incorporate another policy model trained with preference data on tag expansion task to generate augmented tags with high utility. The preference data are collected by tag utility quantification with Shapley value estimation Shapley (1953); Goldshmidt and Horovicz (2024) (detailed in Section 3.4). The tags with high utility are selected as positive while the opposite are set as rejected samples. Initialized from the teacher LLM, the policy model is then optimized and adapted to the tag expansion task with alignment tuning methods like DPO Rafailov et al. (2024). During the tag complexity extension stage, the expanded tags are generated by prompting the optimized policy model with the detailed requirements of semantic alignment and global coherence criteria.
@@ 3.4 Utility Estimation through Shapley Values
We simplify the computation process and estimate the Shapley value based on an existing instruction-response dataset $\mathcal{D}^{t}={(x_{i},y_{i})}_{i=1}^{M}$ such as Alpaca-Cleaned Taori et al. (2023) and Tulu-mixture Lambert et al. (2025). This dataset is different from the instruction data which is to be optimized for better complexity with Tag-Instruct. The tags are extracted from the instructions following the practice in Section 3.2. Following the suggestions in Zhao et al. (2024), we use response length as a proxy for instruction quality $v$, as it effectively captures both complexity and information density. To address the computational complexity of Shapley value calculation, we approximate the tag utility as the average response length of instructions containing that tag:
@@ 3.5 Instruction Reconstruction
The augmented instruction is optimized iteratively, where the output instruction of the $(i-1)^{\mathrm{th}}$ iteration is the input of the $i^{\mathrm{th}}$ iteration with Tag-Instruct.
@@ 4.1 Setup
We conducted experiments using Alpaca-Clean dataset (Taori et al., 2023), from which we randomly sampled 5K instructions to create Alpaca-5k as our initial instruction set. And we use LLaMA-3-8B and LLaMA-3.2-3B as our base models, and Ministral-Instruct-8b (Jiang et al., 2023) as the teacher model. For data pool construction  3.4, we extract tags from instructions following Lu et al. (2023) and utilize instructions from Xu et al. (2024) to estimate tag rewards and train our policy model for tag expansion. Detailed specifications of the datasets, models, and training procedures are provided in Appendix A.
@@
We evaluated our method against several state-of-the-art instruction augmentation methods and popular instruction-tuning datasets. The methodological baselines included Self-Instruct (Wang et al., 2023), Evol-Instruct (Xu et al., 2023), Tree-Instruct (Zhao et al., 2023), Auto-Instruct-Evol (Zeng et al., 2024), and CodecLM (Wang et al., 2024b). For data baselines, we compared against Alpaca-Clean (Taori et al., 2023) and WizardLM-data (Xu et al., 2023). Detailed descriptions of each baseline can be found in Appendix B.
@@ 5.1 Different Crafted Prompts for Semantic Compression
In this part, we compare four crafted prompts for semantic compression which are designed with different strategies. We selected 1,000 instructions from Alpaca-5k and conducted five iterations of complexity enhancement to evaluate four encoding strategies: Basic (simple template), Enhanced (joint intention prioritization and semantic compression), Evolved (intention-focused encoding), and Model-based (pure intention encoding) Lu et al. (2023).
We assessed their performance along two key dimensions: (1) semantic preservation, measured by semantic similarity and ROUGE-L scores between original instruction and reconstructed instruction, and (2) intent extraction capability, which evaluates whether the prompt explicitly requires extracting task intention from instructions. The more detailed results are provided in Appendix H.
As shown in Table 3, different semantic compression strategies exhibit distinct trade-offs between semantic preservation and enhancement capability. The basic template achieves high semantic similarity (0.6536) but limited enhancement (AlpacaEval2.0 LC: 12.9182), while pure model-based encoding Lu et al. (2023) shows stronger enhancement (14.3498) at the cost of lower semantic similarity (0.3600). Through our heuristic design process, we found that jointly encoding both intent and semantic information provides an optimal balance, achieving strong performance on ArenaHard (18.0) while maintaining reasonable semantic preservation.
@@ 5.2 Tag Utility Analysis
To better understand the differences between tags with varying utility values, we categorized the tags into quartiles (Q1-Q4) based on their utility values in ascending order. We then analyzed the number of derived meanings for the tags within each quartile, employing the prompting detailed in Appendix I.1. This prompt is designed according to research on semantic information measurement Kuhn et al. (2023), which systematically probes for as many meanings as possible before merging similar concepts to form the final distinct meanings.
As shown in Figure 4, tags with higher utility values exhibited consistently more derived meanings, indicating they encoded richer semantic information that could be interpreted in more diverse ways. For instance, the tag pytorch from Q4 encompassed multiple semantic meanings, including neural network architectures, deep learning training utilities, and optimization algorithms. This semantic richness enabled diverse instruction generation pathways. A comprehensive analysis of additional representative examples and their semantic interpretations is provided in Appendix J. This finding suggested that the effectiveness of high-utility tags in generating complex and diverse instructions may be attributed to their inherently richer semantic content.
@@ 5.3 The Impact of Different Teacher Model
In this section, we explore the performance of Tag-Instruct with different teacher models. Using the same experimental setup as our main experiments, we compare our approach with baselines by finetuning xx using Qwen2.5-72B-instruct (Qwen et al., 2024) as the teacher model. As shown in Table 4, Tag-Instruct consistently outperforms the baselines, achieving the highest scores on both AlpacaEval2.0 length-controlled tasks (42.76%) and ArenaHard evaluation (41.2%).
These results underscore the broad applicability of our method, surpassing the baseline models across different teacher models. When compared to using Ministral-Instruct-8b as the teacher, Qwen2.5-72B-instruct enhances performance by 23.26 on AlpacaEval2.0 and 19.1 on ArenaHard. This suggests that the superior the performance of the teacher, the higher the quality of the instructions ultimately obtained.
@@ 6 Conclusion
In this paper, we propose Tag-Instruct, a novel framework for enhancing instruction complexity through structured semantic compression and controlled difficulty augmentation. By operating in a compressed tag space rather than raw text, our approach enables more precise and controllable instruction enhancement. Empirical results on AlpacaEval 2 demonstrate that Tag-Instruct significantly outperforms existing instruction augmentation methods. The effectiveness of our approach suggests that structured semantic manipulation is a promising direction for instruction optimization in language models.
@@
Despite the effectiveness of our approach, we identify several limitations. First, our semantic compression process may lose some nuanced information during the discretization of continuous semantic spaces into discrete tags. Second, using response length as a proxy for instruction quality in tag utility estimation might not fully capture all aspects of instruction complexity. Finally, the tag expansion process might introduce computational overhead due to the Shapley value estimation, which could affect the efficiency of large-scale instruction generation.

### v2

@@ Abstract
Tag-Instruct: Controlled Instruction Complexity Enhancement
Through Structure-Based Augmentation
He Zhu1, Zhiwen Ruan2, Junyou Su1, Xingwei He5 Yun Chen4, Wenjia Zhang1,3∗ Guanhua Chen2thanks: *Co-corresponding authors: chengh3@sustech.edu.cn, wenjiazhang@tongji.edu.cn 1 Peking University, 2 Southern University of Science and Technology, 3 Tongji University 4 Shanghai University of Finance and Economics, 5 The University of Hong Kong
@@ 1 Introduction
Figure 1: AlpacaEval 2 Length Control Win Rate. Tag-Instruct achieves the highest win rate among all compared methods. Red bars represent official released models. For other methods, all are fine-tuned on LLaMA3-8B base model, where blue, orange, and gray bars represent methods using Qwen2.5-72B-Instruct as teacher model, methods using Mistral-8B-Instruct as teacher model, and base model respectively. Evaluated using GPT-4o-Mini.
While recent work has made progress in synthetic data generation Zhu et al. (2024); Xu et al. (2024), current instruction datasets still lack sufficient complexity to fully develop model capabilities Xu et al. (2023). Studies show that model performance strongly correlates with instruction difficulty Zhao et al. (2023, 2024), yet existing approaches struggle to effectively control and scale instruction difficulty. Specifically, Xu et al. (2023) and its follow-up works Wang et al. (2024b); Zeng et al. (2024) rely on models’ prior knowledge for self-reflection on augmentation directions, typically using prompts to directly generate “harder” instructions. However, this approach lacks precise control over difficulty progression, as the concept of “harder” remains ambiguous and unquantifiable, making it challenging to systematically guide the augmentation process. While Zhao et al. (2023) attempts to address this by discretizing instructions into semantic trees for manipulation, it still struggles to identify which semantic components are worth augmenting and how to effectively guide the augmentation direction. We name these approaches as Prompt-based Augmentation, which face two fundamental challenges: (1) difficulty in identifying which semantic components are worth augmenting, and (2) lack of principled guidance on which augmentation direction would lead to meaningful complexity increases.
@@
Building on this intuition, we propose Tag-Instruct, a novel Compress & Operation framework for controlled difficulty augmentation. Our framework operates through a three-stage iterative pipeline: (1) semantic encoding that compresses instructions into a compact tag representation while preserving task-relevant information, (2) RL-based tag expansion that explores the tag space to discover new semantic tags that meaningfully increase complexity. (3) instruction synthesis that generates enhanced instructions by conditioning on both the expanded tag set and original instruction. This iterative process enables progressive difficulty increases while maintaining semantic coherence.
Extensive experiments demonstrate the effectiveness of Tag-Instruct across diverse settings. Using Alpaca-5k Taori et al. (2023) as the base dataset, our method enables LLaMA-3-8B to achieve performance comparable to GPT-4-Turbo on standard benchmarks, significantly outperforming existing instruction augmentation approaches. Through detailed ablation studies, we verified that operating in tag space provides superior controllability and makes different instruction synthesis frameworks more stable and effective. [^1] Our code is publicly available at https://github.com/sustech-nlp/Tag-Instruct.
@@ 3.1 Motivation and Overview
The quality of synthesized data correlates to correctness, diversity and complexity Liu et al. (2023b); Zhu et al. (2024). While recent works Wang et al. (2023); Kaur et al. (2025) primarily focus on enhancing instruction diversity, we systematically address instruction complexity enhancement, aligning with previous approaches Xu et al. (2023); Zhao et al. (2023). Following the standard assumption in data synthesis works of having access to a capable instruction-following teacher model, we propose Tag-Instruct for controlled complexity augmentation.
Existing prompt-based augmentation methods Xu et al. (2023); Zhao et al. (2023) operate directly in token space, lacking quantitative control mechanisms and principled guidance for complexity enhancement. To overcome these limitations, Tag-Instruct operates through an Encode → Expansion → Decode paradigm that fundamentally differs from existing token-space methods by compressing instructions into a structured tag space for systematic concept-level operations. For each input instruction $x$, our framework operates through three stages: Stage (1) Encode: compresses instruction $x$ into semantic tags $z_{\text{base}}$ (Section 3.2); Stage (2) Expansion: expands the tag set using RL-guided selection to identify high-utility tags $z_{\text{new}}$, forming $z_{\text{base}^{\prime}}=z_{\text{new}}\cup z_{\text{base}}$ (Section 3.3) guided by Shapley value estimation (Section 3.4); Stage (3) Decode: reconstructs an enhanced instruction $x^{\prime}$ from the expanded tags and original instruction $\{z_{\text{base}}^{\prime},x\}\rightarrow x^{\prime}$ (Section 3.5). This approach is orthogonal to diversity-focused methods and can be integrated with existing synthesis frameworks. Please refer to Figure 2 for the framework overview.
@@ 3.2 Semantic Compression
Given an instruction $x$ from existing synthesized data, the teacher LLM is prompted (see the instruction-to-tag prompt in Appendix E.1) to generate relevant tags $z_{\mathrm{base}}$ of the input instruction. The crafted prompt is expected to explicitly guide the model to extract tags that preserve core semantic meanings with the input. Among these tags, those related to action-oriented concepts and task objectives are emphasized and with higher priority as they are indicative of the input intension. For example, the instruction Create an HTML page that displays a list of books and their authors is compressed into {book_listing, author_info, web_develop}. We discuss and compare different prompt guidelines for semantic compression in Section 5.1.
@@ 3.3 Tag Complexity Expansion
While prompt-based methods can reasonably assess semantic alignment and global coherence through conceptual compatibility and feasibility checks that suit LLM strengths, tag utility quantification requires more sophisticated approaches. Although one could directly prompt the LLM to evaluate tag utility Zheng et al. (2023), this method inadequately models complex tag interactions and provides no principled evaluation framework. Tag-Instruct is proposed to incorporate a policy model trained with preference data on tag expansion task to generate augmented tags with high utility. The preference data are collected by tag utility quantification with Shapley value estimation Shapley (1953); Goldshmidt and Horovicz (2024) (detailed in Section 3.4). The tags with high utility are selected as positive while the opposite are set as rejected samples. Initialized from the teacher LLM, the policy model is then optimized and adapted to the tag expansion task with alignment tuning methods like DPO Rafailov et al. (2024). During the tag complexity extension stage, the expanded tags are generated by prompting the optimized policy model with the detailed requirements of semantic alignment and global coherence criteria.
@@ 3.4 Utility Estimation through Shapley Values
We simplify the computation process and estimate the Shapley value based on an existing instruction-response dataset $\mathcal{D}^{t}={(x_{i},y_{i})}_{i=1}^{M}$ such as Alpaca-Cleaned Taori et al. (2023) and Tulu-mixture Lambert et al. (2025). This dataset is different from the instruction data which is to be optimized for better complexity with Tag-Instruct. The tags are extracted from the instructions following the practice in Section 3.2. Following the suggestions in Zhao et al. (2024) and Shen (2024), we use response length as a proxy for instruction quality $v$, as it effectively captures both complexity and information density. To address the computational complexity of Shapley value calculation, we approximate the tag utility as the average response length of instructions containing that tag:
@@ 3.5 Instruction Reconstruction
The augmented instruction is optimized iteratively, where the output instruction of the $(i-1)^{\mathrm{th}}$ iteration is the input of the $i^{\mathrm{th}}$ iteration with Tag-Instruct. Through this iterative process, each subsequent iteration introduces additional complexity and challenges, as new tags and requirements are incorporated to progressively increase the difficulty level of the instruction.
@@ 4.1 Setup
We conducted experiments using Alpaca-Clean dataset (Taori et al., 2023), from which we randomly sampled 5K instructions to create Alpaca-5k as our initial instruction set. And we use LLaMA-3-8B and LLaMA-3.2-3B as our base models, and Ministral-Instruct-8b (Jiang et al., 2023) as the teacher model. For data pool construction  3.4, we extract tags from instructions following Lu et al. (2023) and utilize instructions from Xu et al. (2024) to estimate tag rewards and train our policy model for tag expansion. Detailed specifications of the datasets, models, and training procedures are provided in Appendix B.
@@
We evaluated our method against several state-of-the-art instruction augmentation methods and popular instruction-tuning datasets. The methodological baselines included Self-Instruct (Wang et al., 2023), Evol-Instruct (Xu et al., 2023), Tree-Instruct (Zhao et al., 2023), Auto-Instruct-Evol (Zeng et al., 2024), and CodecLM (Wang et al., 2024b). For data baselines, we compared against Alpaca-Clean (Taori et al., 2023) and WizardLM-data (Xu et al., 2023). Detailed descriptions of each baseline can be found in Appendix C.
@@ 5.1 Different Crafted Prompts for Semantic Compression
A critical challenge in instruction augmentation is achieving meaningful complexity enhancement while preventing semantic drift from the original instruction. To address this, we compare four prompt strategies for semantic compression, each designed with different priorities. We selected 1,000 instructions from Alpaca-5k and conducted five iterations of complexity enhancement to evaluate four encoding strategies: Basic (simple template), Enhanced (joint intention and semantic compression), Evolved (intention-focused encoding), and Model-based (pure intention encoding) Lu et al. (2023). Detailed prompts for each strategy are provided in Appendix E.1.
We assessed their performance along two key dimensions: (1) semantic preservation, measured by semantic similarity and ROUGE-L scores between original and reconstructed instructions, and (2) intent extraction capability, which evaluates whether the prompt explicitly requires extracting task intention from instructions.
As shown in Table 3, a clear pattern emerges: methods with intent extraction capability (Intent Tag = True) consistently outperform the basic approach, with performance gains inversely correlated with semantic similarity. This reveals a fundamental insight—task-aware compression that prioritizes functional understanding over surface-level similarity yields superior complexity enhancement. The model-based approach achieves the highest AlpacaEval score (14.3498) despite the lowest similarity (0.3600), demonstrating that aggressive semantic abstraction can be beneficial. However, to maintain instruction coherence and prevent excessive semantic drift during iterative augmentation, we adopt the Evolved strategy, which achieves the best ArenaHard performance (18.0) while preserving moderate semantic alignment (0.5373). This strategic trade-off ensures our framework generates complex yet semantically grounded instructions throughout the iterative enhancement process.
@@ 5.2 Tag Utility Analysis
To better understand the differences between tags with varying utility values, we categorized the tags into quartiles (Q1-Q4) based on their utility values in ascending order. We then analyzed the number of derived meanings for the tags within each quartile. The derived meanings measure captures how many distinct semantic interpretations a tag can reasonably support across different contexts. For instance, while database_backup typically refers to a single concept of data preservation, security can encompass authentication mechanisms, encryption protocols, access control systems, and threat detection frameworks. We employed the prompting detailed in Appendix E.2, which is designed according to research on semantic information measurement Kuhn et al. (2023). This prompt systematically probes for multiple meanings before merging similar concepts to form distinct semantic categories.
@@
As shown in Figure 4, tags with higher utility values exhibited consistently more derived meanings, indicating they encoded richer semantic information that could be interpreted in more diverse ways. Tags in Q4 (75-100%) demonstrated the highest semantic complexity with broader meaning distributions, while Q1 tags (0-25%) showed more constrained semantic interpretations. This semantic richness enabled diverse instruction generation pathways. A comprehensive analysis of additional representative examples and their semantic interpretations is provided in Appendix F. This finding suggested that the effectiveness of high-utility tags in generating complex and diverse instructions may be attributed to their inherently richer semantic content.
5.3 Quantitative Control Through High-Utility Tags
We investigate whether our utility-based tag selection provides quantitative control beyond expansion operations. Tag combination, which merges distinct semantic concepts to create diverse instructions, is fundamental to recent synthesis methods Kaur et al. (2024); Didolkar et al. (2024), yet existing approaches lack principled guidance for selection. Our framework transforms this from heuristic sampling to quantitative optimization through Shapley-based utility scores.
To validate this quantitative advantage, we constructed three tag pools of 1000 tags each from Section 3.4 using different selection criteria: High-utility (selecting tags with highest Shapley values), Low-utility (selecting tags with lowest Shapley values), and Random (uniformly sampling tags). For each pool, we randomly sampled two tags to combine and decoded them into instructions using the prompting template in Appendix E.3, generating 5,000 instructions per pool. We then fine-tuned models using these instruction sets under identical configurations to ensure fair comparison.
Table 4 demonstrates the impact of quantitative tag selection. High-utility tags yield 12.3% longer responses (967.52 vs 861.38 tokens) and achieve 36.1% relative improvement on AlpacaEval 2.0 (20.21% vs 14.85%) compared to random selection—despite generating more concise instructions (94.21 vs 99.09 tokens). This efficiency gain (better performance with shorter prompts) validates that our utility scores capture intrinsic semantic value rather than superficial verbosity. Furthermore, the monotonic relationship between tag utility and downstream performance (Low: 11.02% → Random: 14.85% → High: 20.21% on AlpacaEval) establishes that our framework provides quantitative control over instruction quality—a capability absent in existing prompt-based methods. Ultimately, this demonstrates that structured tag-space operations with utility guidance offer principled mechanisms for instruction synthesis, enabling precise optimization rather than stochastic exploration.
Strategy #Inst #Resp AE2.0 AH
Low-utility 85.02 524.87 11.02 15.8
Random 99.09 861.38 14.85 20.9
High-utility 94.21 967.52 20.21 24.6
Table 4: Quantitative impact of tag utility on synthesis quality. High-utility tags consistently outperform random selection across metrics, validating our Shapley-based scoring mechanism.
@@ 5.4 The Impact of Different Teacher Model
In this section, we explore the performance of Tag-Instruct with different teacher models. Using the same experimental setup as our main experiments, we compare our approach with baselines by finetuning LLama3-8B using Qwen2.5-72B-instruct (Qwen et al., 2024) as the teacher model. As shown in Table 5, Tag-Instruct consistently outperforms the baselines, achieving the highest scores on both AlpacaEval2.0 length-controlled tasks (42.76%) and ArenaHard evaluation (41.2%). These results underscore the broad applicability of our method, surpassing the baseline models across different teacher models. When compared to using Ministral-Instruct-8b as the teacher, Qwen2.5-72B-instruct enhances performance by 23.26 on AlpacaEval2.0 and 19.1 on ArenaHard. This significant improvement can be attributed to two key factors: First, stronger models typically exhibit superior instruction-following capabilities, resulting in higher-quality instruction generation with fewer inconsistencies or unsolvable problems. Second, more capable teacher models possess superior reasoning abilities, allowing them to generate higher-quality responses with more sophisticated analytical thinking and deeper insights. This suggests that leveraging more advanced teacher models can substantially enhance both the complexity and quality of the generated instruction dataset.
@@ 6 Conclusion
In this paper, we propose Tag-Instruct, a novel framework for enhancing instruction complexity through structured semantic compression and controlled difficulty augmentation. By operating in a compressed tag space rather than raw text, our approach enables more precise and controllable instruction enhancement. Empirical results on AlpacaEval 2 demonstrate that Tag-Instruct outperforms existing instruction augmentation methods. Our analysis shows that structured semantic manipulation enables controlled progression of instruction difficulty while maintaining coherence across different synthesis frameworks. The effectiveness of our approach suggests that structured semantic manipulation is a promising direction for instruction optimization in language models.
@@
Despite the effectiveness of our approach, we identify several limitations. First, our semantic compression process may lose some nuanced information during the discretization of continuous semantic spaces into discrete tags. Second, using response length as a proxy for instruction quality in tag utility estimation might not fully capture all aspects of instruction complexity. Third, our framework may introduce some noise during the tag expansion process, as we do not incorporate explicit filtering mechanisms to validate the semantic coherence of newly generated tags. Finally, the generalizability of our approach may be limited as our current evaluation focuses primarily on general-domain instructions; future work should explore the effectiveness of Tag-Instruct in specialized domains such as code generation and mathematical reasoning tasks.